OpenStudy (anonymous):

(2c^4) (-c^2) (4c^9) : choices - a) 8c^11 ; b) -8c^15

OpenStudy (lgbasallote):

here's a hint:\[\huge \implies (2 \times -1 \times 4) \times (c^{4 + 2 + 9})\] does that help?

OpenStudy (anonymous):

Omg thankyou!!!!!!

OpenStudy (lgbasallote):

welcome